# https://humanaigc.github.io/emote-portrait-alive/ EMO: Emote Portrait Alive - Generating Expressive Portrait Videos with Audio2Video Diffusion Model under Weak Conditions

### v1

@@ 4.4 Quantitative Comparisons
| Method |
FID |
SyncNet |
F-SIM |
FVD |
E-FID |
| Wav2Lip[18] |
9.38 |
5.79 |
80.34 |
407.93 |
0.693 |
| SadTalker[33] |
10.31 |
4.82 |
84.56 |
214.98 |
0.503 |
| DreamTalk[15] |
58.80 |
3.43 |
67.87 |
619.05 |
2.257 |
| GT |
7.3 |
77.44 |
| Ours | 8.76 |
3.89 |
78.96 | 67.66 | 0.116 |
As show in Table 1, our results demonstrate a substantial advantage in video quality assessment, as evidenced by the lower FVD scores. Additionally, our method outperforms others in terms of individual frame quality, as indicated by improved FID scores. Despite not achieving the highest scores on the SyncNet metric, our approach excels in generating lively facial expressions as shown by E-FID.
5 Limitation
There are some limitations for our method. First, it is more time-consuming compared to methods that do not rely on diffusion models. Second, since we do not use any explicit control signals to control the character’s motion, it may result in the inadvertent generation of other body parts, such as hands, leading to artifacts in the video. One potential solution to this issue is to employ control signals specifically for the body parts.

### v2

@@ 4.4 Quantitative Comparisons
| Method | FID | SyncNet | F-SIM | FVD | E-FID |
| Wav2Lip[23] | 9.38/31.70 | 5.79/4.14 | 80.34/78.87 | 407.93/487.00 | 0.693/0.652 |
| SadTalker[40] | 10.31/31.37 | 4.82/2.90 | 84.56/81.86 | 214.98/418.19 | 0.503/0.539 |
| DreamTalk[19] | 58.80/88.21 | 3.43/1.29 | 67.87/56.38 | 619.05/584.63 | 2.257/3.548 |
| MakeItTalk[42] | 21.73/39.86 | 2.85/1.64 | 76.91/60.12 | 350.96/340.55 | 1.072/0.997 |
| GT | - | 7.3/2.69 | 77.44/72.64 | - | - |
| w/o 250h data | 10.80/- | 5.02/- | 79.55/- | 102.78/- | 0.215/- |
| Ours | 8.76/17.33 | 3.89/2.74 | 78.96/77.16 | 67.66/192.77 | 0.116/0.187 |
Figure 3 illustrates that the internet dataset encompasses a wider variety of facial expressions and an extensive range of head movements, coupled with diverse poses of the reference character. Such variability has the potential to negatively impact performance metrics, as shown in Table 1. our results demonstrate a substantial advantage in video quality assessment, as evidenced by the lower FVD scores. Additionally, our method outperforms others in terms of individual frame quality, as indicated by improved FID scores. Wav2Lip has the highest SyncNet confidence score by training with SyncNet as a dicriminator[19], despite not achieving the highest scores on the SyncNet metric, our approach excels in generating lively facial expressions as shown by E-FID. Furthermore, we also investigated the benefit of our module and the impact of the 250 hours dataset by training our model solely on publicly available datasets including VFHQ, HDTF, and CELEB-V [43]. Even in the absence of the 250-hour dataset, our model still exhibits exceptional performance, particularly in terms of FVD and E-FID. While the further improvements on these metrics caused by the collected dataset demonstrates that the extra data could contributes to enhancing video content dynamics and generating a wider variety of expressions.
4.5 Ablation Studies
4.5.1 the Impact of the Speed Layer.
The speed layers are designed to ensure the consistency of the head motion frequency between the contiguous generated video clips. During the inference, we assign a constant speed value for every frame. To measure the effect of the speed layers, we generate video clips with different assigned speed value on the HDTF dataset. As shown in Table 2, "No Speed" indicates the results generated by the model without the speed layers. "Velocity Variance" represents the average variance of the velocities across individual video sequences, reflecting the consistency of rotational speed within each clip. "Variance of Mean Velocities (VMV)" indicates the variance of the mean velocities across different clips, providing a measure of the variability in head rotation speeds from one clip to the others. Incorporating speed layers significantly enhances the stability of head motion, as evidenced by the reduced "Velocity Variance" and "VMV" in comparison to the baseline "No Speed" condition. These metrics demonstrate that the model with speed layers yields more consistent head rotation velocities within clips, respectively. Furthermore, the "Mean Velocity" metric substantiates that the pre-defined speed value affect the actual velocity level, confirming the efficacy of the speed layers in modulating the synthesized head motion dynamics. In our approach, speech-driven cases are assigned speeds from 0.1 to 1.0, while singing scenarios might use higher settings (1.0-1.3) for faster, song-congruent head motions. Speeds exceeding 1.5 could lead to unnaturally rapid and jittery movements.
| No Speed | 0.1 | 0.7 | 1.3 | 1.9 | |
| Mean Velocity | 1.365 | 0.878 | 1.001 | 1.162 | 1.246 |
| Velocity Variance | 1.002 | 0.357 | 0.454 | 0.550 | 0.657 |
| VMV | 0.134 | 0.046 | 0.054 | 0.057 | 0.058 |
4.5.2 the Control Effect of the Face Locator.
Face locator take the face region as input, and delineates the permissible domain for facial movements, thereby influencing the range of head motion. This is illustrated in Figure 7, where the character exhibits minimal head movement upon receiving an appropriately sized facial region as input. Conversely, a broader input region permits the character to exhibit more head swings during speech, and an input region with increased height facilitates nodding gestures. Furthermore, inputting a uniform white mask does not provide specific guidance, allowing for facial generation in arbitrary locations. The purple bboxes could extend beyond the prescribed white face region, evidencing that the face locator exerts only a weak condition on head movements, permitting a degree of motion that transcends its indicated boundaries.
5 Conclusion
In this work, we introduced EMO, a framework that advances the generation of expressive talking head videos by leveraging audio input to drive the animation process. By eschewing the traditional reliance on intermediate signals, EMO capitalizes on weakly conditioned audio2video diffusion to generate lifelike portraits. Our method demonstrates notable improvements in capturing the nuances of human expressiveness and maintaining consistent identity across video frames. The results outperform existing state-of-the-art methods, confirming the potential of EMO as a powerful tool for creating rich, audio-driven video content.